A couple has just purchased a home for $425,300.00. They will pay 20% down in cash, and finance the remaining balance. The mortgage broker has gotten them a mortgage rate of 3.36% APR with monthly compounding. The mortgage has a term of 30 years.

How much interest is paid in the first year?
Answer #1

w N A B с D E F 1. Payment Beg amount Payment Int. portion Pricipal portion Principal bal. 2 1 $340,240.00 $1,501.36 $ 952.67

Calc:-

A B С D E F 1 Payment Beg amount Payment Int. portion Pricipal portion Principal bal. 2 1 =425300*80% =PMT(3.36%/12,30*12,-42